# Stability of Thin Shell and Wormhole Configurations: Schwarzschild, Schwarzschild - (Anti-) de Sitter, and FLRW Spacetimes

### v1

@@ V Results
In the following, we will turn our attention to the plotting of regions of stability in three-dimensional parameter space. Overall qualitative features of the observed stability regions will be discussed.
@@ V.1 Plotting
We can now view the parameter space by plotting $P^{\prime}/\sigma^{\prime}$ as a function of $\alpha$ and $\beta$. Since $m^{+}$ is not a function of $R$, when plugging into equations (22) and (23) it will cancel and have no effect on the final stability surface. Following a similar approach as [6], we shall let $|C^{\pm}|=|\Lambda^{\pm}m^{+2}|\sim 10^{-3}$. This assumption gives a high value for this product, but is useful for qualitative plotting.
@@
since $01-\beta>$ 0, and
$[C]>\frac{-2C_{\rm max}^{\frac{3}{2}}}{\sqrt{3}}(1-\beta).$ (62)
@@
For the junctions between FLRW and Schwarzschild, we consider $\alpha=\chi/m$ and $C^{\pm}=\Lambda^{\pm}m^{2}$. Note that $\beta=m^{-}/m=0$ in this case as $m^{-}=0$. Thus, for plotting, we will use the $\alpha$ and $a(t)$ axes.
We choose $C^{+}=\pm 10^{-3}$ for Schwarzschild - de Sitter or Schwarzschild - anti-de Sitter and $C^{-}=+10^{-3}$ always. Thus, $[C]\leq 0$.
@@
Again, since $m$ is constant with respect to $\chi$, it will cancel when plugged into equations (49) and (50). However, the contribution from curvature $k$ is affected by the $m^{2}$ factor which cannot be ignored. A value for this $m$ must be chosen when plotting $k\neq 0$.
(a) Black hole
(b) Wormhole
Figure 12: FLRW k=-1 – Schwarzschild Junction: There is no horizon present at high $\chi$ for neither the black hole or the wormhole. For high $\chi$ the value of the surface approaches 0 instead of $\sim-0.5$. Stability regions are below A and above B. Surface B of wormhole graph decreases parabolically at higher $\chi$ than plotted. Stability regions are above A and below B.
@@
However, due the fact that $[C]\leq 0$, equation (64) is never satisfied when $k\neq+1$.
@@
$[C]>\frac{D}{a^{3}},$ (66)
which is never satisfied for $[C]\leq 0$.
When using equation (59) we assume $\varepsilon_{\rm max}=\varepsilon^{-}=3km^{2}$ as $D>0$ and $[C]<0$ cause $\varepsilon^{-}>\varepsilon^{+}$. This condition will only be relevant if $\varepsilon^{-}>0$ which implies the $k=+1$ case. So,
$[C]>\frac{D-6m^{3}}{a^{3}},$ (67)
which implies if $D-6m^{3}>0$ there can be no asymptote at any $a(t)$ for $[C]\leq 0$. Additionally, $\alpha_{\rm dS}\approx\frac{1}{m}$ when $k=+1$. Such a horizon is nonexistent for $k\neq+1$.
Plotted examples of parameter spaces for Schwarzschild and Schwarzschild - (anti-) de Sitter junctions are given in figures (1, 2, 3, 4, 5, 6, 7). Figures (8, 9, 10, 11, 12) detail parameter spaces of FLRW junctions with Schwarzschild and Schwarzschild - (anti-) de Sitter.
Throughout the plots, there are typically two separate regions of stability. Region A will refer to the stability region at lower $R$ or $\chi$ than the $\Upsilon=0$ asymptote or the $M=0$ stability flip. Region B shall refer to the higher $R$ or $\chi$ region [6].
@@ V.2 Discussion
From the results seen in the figures throughout and summarized in Tables 1 and 2, we can see that the predictions made in Section III and IV are met. All (FLRW included) wormhole junctions possess an asymptote at $\Upsilon=0$. The shape and location of the asymptote can vary, but in all cases it divides the parameter space into two distinct stability regions. Stability region A is at lower $R$ or $\chi$ and is bound by a minimum $\frac{P^{\prime}}{\sigma^{\prime}}>0$. Stability region B is at higher $R$ and is bound by a maximum $\frac{P^{\prime}}{\sigma^{\prime}}<0$. In all cases observed these stability regions do not intersect with the Causal Region, $10\leq\frac{P^{\prime}}{\sigma^{\prime}}<$ 1.
All (FLRW included) black hole junctions either possess an asymptote or an $M=0$ stability flip with the exception of the simple Schwarzschild – Schwarzschild and Schwarzschild - anti-de Sitter – Schwarzschild - anti-de Sitter. Of those that have some form of stability flip, (either due to $M=0$ or $\Upsilon=0$) there are two separate stability regions. Region A is at low $R$ or $\chi$ and is defined by a maximum $\frac{P^{\prime}}{\sigma^{\prime}}$. Region B is at higher R and is defined by a minimum $\frac{P^{\prime}}{\sigma^{\prime}}$. The maximum is above $\frac{P^{\prime}}{\sigma^{\prime}}=0$, causing an intersection of the stability regions with the Causal region. However, such an intersection does not exist for every combination of $\alpha$ and $\beta$ or $\alpha$ and $a(t)$.
@@ VI Conclusion
We performed a stability analysis of thin shell wormholes and black holes constructed from combinations of Schwarzschild, Schwarzschild - de Sitter, Schwarzschild - anti-de Sitter, and Friedmann-Lemaître-Robertson-Walker (FLRW) spacetimes. We aimed at providing a taxonomy of these combinations while emphasizing some common landscape characteristics with the localization of stability and causality regions.
As previously stated, we found no parameter combinations which yield a stable wormhole when restricting $10\leq\frac{P^{\prime}}{\sigma^{\prime}}<$ 1. The boundaries of the stability regions consistently lie far outside the causal region, either at very high values or at negative ones. This conclusion has been echoed in previous works see e.g., [3, 6, 7, 32]. Our investigation of junctions containing the FLRW spacetime have yielded the same results.
One may interpret $P^{\prime}/\sigma^{\prime}$ as the square of the sound speed of perturbation, in which case it would be necessary to restrict its values to the causal region where the sound speed could not exceed the speed of light and be restricted to the reals. In this case, one would tend to deny the stability of the considered wormholes in the physical universe. However, as it was cautioned in [3, 7, 32], it may not be fully justified to entirely rule stability out. Indeed, the geometry of a wormhole already requires a violation of the null energy condition where the surface energy density $\sigma<0$ [3]. Such unusual conditions could allow for $P^{\prime}/\sigma^{\prime}$ to take on values outside of the causal range. It has been noted that negative $P^{\prime}/\sigma^{\prime}$ values can occur with the Casimir Effect and the False Vacuum [3]. Unfortunately, until a detailed model of exotic matter is formulated, it remains unclear if such values should be excluded from the discussion or not. It is also of note that possible stability in the causal region has been noted in rotating (BTZ metric) wormholes for sufficient angular momentum [20]. Thus, these wormholes do not require a relaxation of the $10<P^{\prime}/\sigma^{\prime}<$ 1 condition to be stable.

### v2

@@ V Results
In the following, we turn our attention to the plotting of regions of stability in three-dimensional parameter space. Overall qualitative features of the observed stability regions are discussed.
@@ V.1 Plotting
We can now view the parameter space by plotting $P^{\prime}/\sigma^{\prime}$ as a function of $\alpha$ and $\beta$. Since $m^{+}$ is not a function of $R$, when plugging into equations (22) and (23), it cancels and have no effect on the final stability surface. Following a similar approach as [8], we choose $|C^{\pm}|=|\Lambda^{\pm}m^{+2}|\sim 10^{-3}$. This assumption gives a high value for this product, but is useful for qualitative plotting.
@@
since $1-\beta>0$, and
$C]>\frac{-2C_{\rm max}^{\frac{3}{2}}}{\sqrt{3}}(1-\beta).$ (62)
@@
For the junctions between FLRW and Schwarzschild, we consider $\alpha=\chi/m$ and $C^{\pm}=\Lambda^{\pm}m^{2}$. Note that $\beta=m^{-}/m=0$ in this case as $m^{-}=0$.
We choose $C^{+}=\pm 10^{-3}$ for Schwarzschild - de Sitter or Schwarzschild - anti-de Sitter and $C^{-}=+10^{-3}$ always. Thus, $C]\leq 0$.
@@
Again, since $m$ is constant with respect to $\chi$, it cancels when plugged into equations (49) and (50). However, the contribution from curvature $k$ is affected by the $m^{2}$ factor which cannot be ignored. A value for this $m$ must be chosen when plotting $k\neq 0$.
Figure 12: FLRW (k=-1) – Schwarzschild Junction: There is no horizon present at high $\chi$ for either the black hole or the wormhole. As in figure (7), the stability region of the wormhole approaches the lower bound of the Causal Region as $\alpha\rightarrow\infty$. For 3-D plot, see figure (18).
@@
However, due the fact that $C]\leq 0$, equation (64) is never satisfied when $k\neq+1$.
@@
$C]>\frac{D}{a^{3}},$ (66)
which is never satisfied for $C]\leq 0$.
When using equation (59) we assume $\varepsilon_{\rm max}=\varepsilon^{-}=3km^{2}$ as $D>0$ and $[C]<0$ cause $\varepsilon^{-}>\varepsilon^{+}$. This condition is only relevant if $\varepsilon^{-}>0$ which implies the $k=+1$ case. So,
$C]>\frac{D-6m^{3}}{a^{3}},$ (67)
which implies if $D-6m^{3}>0$ there can be no asymptote at any $a(t)$ for $C]\leq 0$. Additionally, $\alpha_{\rm dS}\approx\frac{1}{m}$ when $k=+1$. Such a horizon is nonexistent for $k\neq+1$.
Plotted examples of parameter spaces for Schwarzschild and Schwarzschild - (anti-) de Sitter junctions are given in figures (1-7). Figures (8-12) detail parameter spaces of FLRW junctions with Schwarzschild and Schwarzschild - (anti-) de Sitter. Three dimension plots can be found in Appendix A for select junctions.
Type Asymptote at $\Upsilon=0$ Limit as $R\rightarrow\infty$ Stability in Causal Region Figure
Sch – Sch No Convergent Yes 1
Sch – Sch-deSit Yes Divergent Yes 4
Sch – Sch-Anti deSit No Convergent Yes 5
Sch-deSit – Sch No Divergent Yes 2
Sch-deSit – Sch-deSit Yes Divergent Yes 6
Sch-deSit – Sch-Anti deSit No Divergent Yes -
Sch-Anti deSit – Sch Yes Convergent Yes 3
Sch -Anti deSit – Sch-deSit Yes Divergent Yes -
Sch-Anti deSit – Sch-Anti deSit No Convergent Yes 7
FLRW (k=0) – Sch No Convergent Yes 8
FLRW (k=0) – Sch-deSit No Convergent Yes 9
FLRW (k=0) – Sch-Anti deSit No Convergent Yes 10
FLRW (k=+1) – Sch No Divergent Yes 11
FLRW (k=-1) – Sch No Convergent Yes 12
Table 1: Summary of the notable features of the parameter spaces that explain the stability of a black hole in the Causal Region.
Type Asymptote at $\Upsilon=0$ Limit as $R\rightarrow\infty$ Stability in Causal Region Figure
Sch – Sch Yes Convergent No 1
Sch – Sch-deSit Yes Divergent No 4
Sch – Sch-Anti deSit Yes Convergent No 5
Sch-deSit – Sch Yes Divergent No 2
Sch-deSit – Sch-deSit Yes Divergent No 6
Sch-deSit – Sch-Anti deSit Yes Divergent No -
Sch-Anti deSit – Sch Yes Convergent No 3
Sch -Anti deSit – Sch-deSit Yes Divergent No -
Sch-Anti deSit – Sch-Anti deSit Yes Convergent No 7
FLRW (k=0) – Sch Yes Convergent No 8
FLRW (k=0) – Sch-deSit Yes Convergent No 9
FLRW (k=0) – Sch-Anti deSit Yes Convergent No 10
FLRW (k=+1) – FLRW Yes Divergent No 11
FLRW (k=-1) – FLRW Yes Convergent No 12
Table 2: Summary of the notable features of the parameter spaces that explain the stability of a wormhole in the Causal Region.
@@ V.2 Discussion
From the results seen in the figures throughout and summarized in Tables 1 and 2, we can see that the predictions made in Section III and IV are met. All (FLRW included) wormhole junctions possess an asymptote at $\Upsilon=0$. The location of the asymptote can vary, but in all cases it divides the parameter space into two disjoint stability regions. The stability region at lower $R$ or $\chi$ is bound by a minimum $\frac{P^{\prime}}{\sigma^{\prime}}>0$. The stability region at higher $R$ and is bound by a maximum $\frac{P^{\prime}}{\sigma^{\prime}}<0$. In all cases observed these stability regions do not intersect with the Causal Region, $0\leq\frac{P^{\prime}}{\sigma^{\prime}}<1$.
All (FLRW included) black hole junctions either possess an asymptote or an $M=0$ stability flip with the exception of the simple Schwarzschild – Schwarzschild and Schwarzschild - anti-de Sitter – Schwarzschild - anti-de Sitter. Of those that have some form of stability flip, (either due to $M=0$ or $\Upsilon=0$) the stability is defined by two regions. The region at low $R$ or $\chi$ and is defined by a maximum $\frac{P^{\prime}}{\sigma^{\prime}}$. Region B is at higher R and is defined by a minimum $\frac{P^{\prime}}{\sigma^{\prime}}$. The maximum is above $\frac{P^{\prime}}{\sigma^{\prime}}=0$, causing an intersection of the stability regions with the Causal region. However, such an intersection does not exist for every combination of $\alpha$ and $\beta$ or $\alpha$ and $a(t)$.
@@ VI Conclusion
We performed a stability analysis of thin shell wormholes and black holes constructed from combinations of Schwarzschild, Schwarzschild - de Sitter, Schwarzschild - anti-de Sitter, and Friedmann-Lemaître-Robertson-Walker (FLRW) spacetimes. We provide a taxonomy of these combinations while emphasizing some common landscape characteristics with the localization of stability and causality regions.
We found no parameter combinations which yield a stable wormhole when restricting $0\leq\frac{P^{\prime}}{\sigma^{\prime}}<1$. The boundaries of the stability regions consistently lie outside the causal region, either at very high values or at negative ones. This conclusion was echoed in previous works [24, 8, 18, 20]. Our investigation of junctions containing the FLRW spacetime have yielded the same results.
One may interpret $P^{\prime}/\sigma^{\prime}$ as the square of the sound speed of perturbation, in which case it would be necessary to restrict its values to the causal region where the sound speed could not exceed the speed of light and be restricted to the reals. In this case, one would tend to deny the stability of the considered wormholes in the physical universe. However, as it was cautioned in [24, 18, 20], it may not be fully justified to entirely rule stability out. Indeed, the geometry of a wormhole already requires a violation of the null energy condition [24]. Such unusual conditions could allow for $P^{\prime}/\sigma^{\prime}$ to take on values outside of the causal range. It has been noted that negative $P^{\prime}/\sigma^{\prime}$ values can occur with the Casimir Effect and the False Vacuum [24]. Unfortunately, until a detailed model of exotic matter is formulated, it remains unclear if such values should be excluded from the discussion or not. It is also of note that possible stability in the causal region has been noted in rotating (BTZ metric) wormholes for sufficient angular momentum [30]. Thus, these wormholes do not require a relaxation of the $0<P^{\prime}/\sigma^{\prime}<1$ condition to be stable.
@@ Acknowledgements
Appendix A 3-D Plots
In this section, the parameter spaces of a few junctions are plotted in three dimensions for a more complete visualization. For the junctions with Schwarzschild - (anti) de Sitter (figures 13-15) $P^{\prime}/\sigma^{\prime}$ is plotted against the $\alpha$ and $\beta$ axes and for junctions with FLRW (figures 16-18), it is plotted against the $\alpha$ and $a(t)$ axes.
(a) Black hole
(b) Wormhole
Figure 13: Schwarzschild – Schwarzschild Junction: For the black hole, $M\neq 0$ everywhere and there is no stability flip. Stability partially intersects the Causal Region. For the wormhole there is an asymptote present and stability is separated into two disjoint regions, neither of which intersect the Causal Region. For 2-D plot, see figure (1).
(a) Black hole
(b) Wormhole
Figure 14: Schwarzschild - de Sitter – Schwarzschild Junction: For the black hole, there are points where $M=0$ and a stability flip is present without an asymptote. This occurs at the vertical plane. A Schwarzschild - de Sitter – Schwarzschild junction implies [C]=0.001[C]=-0.001 which does not fulfill the asymptote condition for any beta. For the black hole plot, the de Sitter Horizon can be seen by the divergence around $\alpha\approx 54$, though the view window of the wormhole plot is too restricted for this feature to be visible. For 2-D plot, see figure (2).
(a) Black hole
(b) Wormhole
Figure 15: Schwarzschild - anti-de Sitter – Schwarzschild Junction: Black hole has an asymptote as $[C]=0.001$ which does fulfill the asymptote condition. A de Sitter horizon is not present as $\Lambda^{\pm}\leq 0$. For 2-D plot, see figure (3).
(a) Black hole
(b) Wormhole
Figure 16: FLRW (k=0) – Schwarzschild Junction: The black hole and wormhole graphs are similar to previous examples though does not posses a de Sitter Horizon, despite there existing a $\Lambda>0$. It is worth noting that the radius of the event horizon decreases below $\alpha=2$ at low $a$ due to a high value of $H^{2}$. For 2-D plot, see figure (8).
(a) Black hole
(b) Wormhole
Figure 17: FLRW (k=+1) – Schwarzschild Junction: For $k\neq 0$ we choose $m=0.02$. It is important to note that $k>0$ gives the non physical result of $H^{2}<0$ for $0.420\lesssim a(t)\lesssim 0.823$. A horizon caused by the positive curvature exists at high $\chi$ for both black hole and wormhole. At larger $m$ values the asymptote condition can be fulfilled but commonly this occurs only when $H^{2}<0$. For 3-D plot, see figure (11).
(a) Black hole
(b) Wormhole
Figure 18: FLRW (k=-1) – Schwarzschild Junction: There is no horizon present at high $\chi$ for either the black hole or the wormhole. As in figure (7), the stability region of the wormhole approaches the lower bound of the Causal Region as $\alpha\rightarrow\infty$. For 2-D plot, see figure (12).